OpenStudy (anonymous):

solve the system using an equivalent system {-2x-y=0 {3x-y=10 A. (4,-2) B. (-2,4) C. (2,-4) D. (-4,2)

OpenStudy (senkai):

You would plug in all of the plots into the equations... So for example, (4,-2). Plug it in, so the equation would be: -2(4) - (-2) = 0 and 3(4) - (-2) = 10 Simplified into: -6 + 2 = 0 and 12 + 2 = 10 These are both wrong, so you know the answer is not A. Try doing this for B, C, and D, until you find points that make both equations correct.

OpenStudy (anonymous):

okay thank you:)